OpenStudy (anonymous):

Solve the equation. 2 (x+4) = 30

OpenStudy (anonymous):

2(x+4)=30 2x+8=30 -8 -8 ---------- 2x=22 2x/2=22/2 x=11

OpenStudy (markrijckenberg):

It is a little easier to first divide LHS and RHS by 2. Then isolate x.

OpenStudy (anonymous):

x+4 = 15, x = 11